# disposition

### v2

@@ 3 Scalarization in Vector Optimization
  • $3^{\circ}$.
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $s=\varphi^{G}_{a-K,k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
Proof 3.15.
The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is an immediate consequence of Lemma 3.8 ($4^{\circ}$) taking into account that $\operatorname{int}\,K\subseteq K\setminus\{0\}$, or equivalently $K\neq Y$ (since (3.1) has a solution), while the proof of the implication $1^{\circ}\Longrightarrow 2^{\circ}$ is similar to the proof of the corresponding implication in Proposition 3.11.
If $K$ is closed, the equivalence of $2^{\circ}$ and $3^{\circ}$ follows immediately from Lemma 3.6.
Remark 3.16.
Assuming that $K$ has a normlike-base $B_{K}(\psi)$ (which means that the seminorm $\psi:Y\to\mathbb{R}$ is positive on $K\setminus\{0\}$) and $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq a+K\setminus\ell(K)$ (respectively, $f[{\mathrm{W}Eff}(\Omega\mid f,K)]\subseteq a+\operatorname{int}\,K$) for some $a\in Y$, in Proposition 3.11 (respectively, Proposition 3.14), one may take $\mathcal{P}_{a}:=\{a\}$, $\mathcal{P}_{k}:=B_{\{0\}\cup(K\setminus\ell(K))}(\psi)$ (respectively, $\mathcal{P}_{k}:=B_{\{0\}\cup\operatorname{int}\,K}(\psi)$ if $0\notin\operatorname{int}\,K$, or equivalently, $K\neq Y$) and $\mathcal{R}:=\mathbb{P}$.
The following proposition provides possible choices for the sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq K\setminus\ell(K)$ and $\mathcal{R}\subseteq\mathbb{R}$ to ensure $f[D]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$ with $D={\mathrm{E}ff}(\Omega\mid f,K)$, $D={\mathrm{W}Eff}(\Omega\mid f,K)$ and $D=\Omega$, respectively, in Propositions 3.11 and 3.14.
Proposition 3.17.
@@
Proof 3.18.
$1^{\circ}$. Let us verify the equality $H_{y^{*},a}^{>}=H_{y^{*},a}^{=}+\mathbb{P}\cdot k$. Since $a\in H_{y^{*},a}^{=}$, $y^{*}(k)>0$ and $\mathbb{P}\cdot k\subseteq K\setminus\{0\}$, we conclude $H_{y^{*},a}^{=}+\mathbb{P}\cdot k\subseteq H_{y^{*},a}^{>}$. Take some $y\in H_{y^{*},a}^{>}$. Noting that $y-tk\in H_{y^{*},a}^{=}$ for some $t>0$, we derive $y=y-tk+tk\in H_{y^{*},a}^{=}+\mathbb{P}\cdot k$, which shows the reverse inclusion. Applying the result for the cone $-K$, the functional $-y^{*}$ and the direction $-k$ (in the role of $K$, $y^{*}$ and $k$), we get immediately $H_{y^{*},a}^{<}=H_{(-y^{*}),a}^{>}=H_{(-y^{*}),a}^{=}+\mathbb{P}\cdot(-k)=H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
$2^{\circ}$. Assume that $y^{*}\in K^{\#}$, which means that $y^{*}(x)>0$ for all $x\in K\setminus\{0\}$. Then, using $1^{\circ}$ we easily derive $a+K\setminus\{0\}\subseteq H_{y^{*},a}^{>}=H_{y^{*},a}^{=}+\mathbb{P}\cdot k$. Applying the result for $-K$, $-y^{*}$ and $-k$ (in the role of $K$, $y^{*}$ and $k$), we obtain $a-K\setminus\{0\}\subseteq H_{-y^{*},a}^{=}-\mathbb{P}\cdot k=H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
$3^{\circ}$. Define $C:=({\mathrm{i}nt}\,K)\cup\{0\}$. Note that ${\mathrm{i}nt}\,K=C\setminus\{0\}$ and $K^{+}\setminus\{0\}=\{x^{*}\in Y^{*}\mid\forall x\in{\mathrm{i}nt}\,K:x^{*}(x)>0\}=C^{\#}$ by [50, Th. 4.8, Cor. 4.9] (if $K$ is a solid, convex cone). Applying assertion $2^{\circ}$ for $C$ in the role of $K$ we get the result.
$4^{\circ}$. The first equality is obvious, while the second equality follows by $1^{\circ}$.
Remark 3.19.
@@
Let the assumptions of Proposition 3.17 be fulfilled. Eichfelder’s adaptive scalarization method in multiobjective optimization proposed in [12, 13] is based on parameterized scalar optimization problems of type (3.1) and (3.2), respectively, with different choices of the reference point $a\in H_{y^{*},a}^{=}$ and an a-priori fixed direction $k\in K\setminus\{0\}$. Note that Proposition 3.11 generalizes a result by Eichfelder [13, Th. 3.2] taking into account that $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq Y=\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$ for $\mathcal{P}_{a}:=H_{y^{*},a}^{=}$, $\mathcal{P}_{k}:=\{k\}$ and $\mathcal{R}:=\mathbb{R}$ by Proposition 3.17 ($4^{\circ}$).
4 Nonlinear Cone Separation Conditions

### v3

@@ 3 Scalarization in Vector Optimization
  • $3^{\circ}$
    There are some $a\in\mathcal{P}$, $k\in\mathcal{T}$ and $s\in\mathcal{R}$ so that $s=\varphi^{G}_{a-K,k}(f(\bar{x}))={\rm min}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
Proof.
The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is an immediate consequence of Lemma 8 ($4^{\circ}$) taking into account that $\operatorname{int}\,K\subseteq K\setminus\{0\}$, or equivalently $K\neq Y$ (since (7) has a solution), while the proof of the implication $1^{\circ}\Longrightarrow 2^{\circ}$ is similar to the proof of the corresponding implication in Proposition 9 by using Lemma 8 ($2^{\circ}$).
If $K$ is closed, the equivalence of $2^{\circ}$ and $3^{\circ}$ follows immediately from Lemma 7. ∎
Remark 8.
Assuming that $K$ has a normlike-base $B_{K}(\psi)$ (which means that the seminorm $\psi:Y\to\mathbb{R}$ is positive on $K\setminus\{0\}$) and $f[{\rm Eff}(\Omega\mid f,K)]\subseteq a+K\setminus\ell(K)$ (respectively, $f[{\rm WEff}(\Omega\mid f,K)]\subseteq a+\operatorname{int}\,K$) for some $a\in Y$, in Proposition 9 (respectively, Proposition 10), one may take $\mathcal{P}:=\{a\}$, $\mathcal{T}:=B_{\{0\}\cup(K\setminus\ell(K))}(\psi)$ (respectively, $\mathcal{T}:=B_{\{0\}\cup\operatorname{int}\,K}(\psi)$ if $0\notin\operatorname{int}\,K$, or equivalently, $K\neq Y$) and $\mathcal{R}:=\mathbb{P}$.
The following proposition provides possible choices for the sets $\mathcal{P}\subseteq Y$, $\mathcal{T}\subseteq K\setminus\ell(K)$ and $\mathcal{R}\subseteq\mathbb{R}$ to ensure $f[D]\subseteq\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$ with $D={\rm Eff}(\Omega\mid f,K)$, $D={\rm WEff}(\Omega\mid f,K)$ and $D=\Omega$, respectively, in Propositions 9 and 10.
Proposition 11.
@@
Proof.
The proofs for $1^{\circ}$ and $2^{\circ}$ are straightforward, while $3^{\circ}$ and $4^{\circ}$ follow easily by using $1^{\circ}$ and $2^{\circ}$, taking into account that, for a nontrivial, solid, convex cone $K$ and $C:=({\rm int}\,K)\cup\{0\}$ we have ${\rm int}\,K=C\setminus\{0\}$ and $K^{+}\setminus\{0\}=\{x^{*}\in Y^{*}\mid\forall x\in{\rm int}\,K:x^{*}(x)>0\}=C^{\#}$ (see also [48, Th. 4.8, Cor. 4.9]). ∎
Remark 9.
@@
Let the assumptions of Proposition 11 be fulfilled. Eichfelder’s adaptive scalarization method in multiobjective optimization proposed in [4, 5] is based on parameterized scalar optimization problems of type (7) and (8), respectively, with different choices of the reference point $a\in H_{y^{*},a}^{=}$ and an a-priori fixed direction $k\in K\setminus\{0\}$. Note that Proposition 9 generalizes a result by Eichfelder [5, Th. 3.2] taking into account that $f[{\rm Eff}(\Omega\mid f,K)]\subseteq Y=\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$ for $\mathcal{P}:=H_{y^{*},a}^{=}$, $\mathcal{T}:=\{k\}$ and $\mathcal{R}:=\mathbb{R}$ by Proposition 11 ($4^{\circ}$).
4 Bishop-Phelps Type Scalarization in Vector Optimization
Assume that the real topological-linear space $Y$ is endowed with a nontrivial cone $K\subseteq Y$. Consider a sublinear functional $\varphi:Y\to\mathbb{R}$ with $\varphi(0)=0$. As highlighted in Section 1, for any given point $a\in Y$ and any direction $k\in Y\setminus\{0\}$, the following scalarizations of the vector problem ($P$) will play a key role in our studies:
$\begin{cases}\varphi(f(x)-a)\to\min\\ x\in\Omega\end{cases}$ ($P_{\varphi}^{a}$)
and
$\begin{cases}\lambda\to\min\\ \varphi(f(x)-a-\lambda k)\leq 0,\\ (x,\lambda)\in\Omega\times\mathbb{R}.\end{cases}$ ($P_{\varphi}^{a,k}$)
When dealing with the scalar problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$), we will impose that $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing for some $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$. Thus, the Bishop-Phelps type cone $C_{\psi}(x^{*},\alpha)$ plays an important role in our scalarization approach due to the cone representing property of $\varphi$. This is why our method is called Bishop-Phelps type scalarization in vector optimization.
Remark 10.
It is known that several scalarizing functions in vector optimization (e.g., in the sense of Gerstewitz [26, 27]; Hiriart-Urruty [53] / Zaffaroni [11, Prop. 3.2]) admit cone representation properties. Such functions can therefore be used well in our approach. Let us describe two types of scalarizing functions in more detail:
  • The nonlinear scalarizing function $\varphi^{G}_{-C_{\psi}(x^{*},\alpha),k}$ (i.e., the function defined in (9) with $a=0$ and with $C_{\psi}(x^{*},\alpha)$ in the role of $K$) is sublinear and $-C_{\psi}(x^{*},\alpha)$ representing (respectively, strictly $-C_{\psi}(x^{*},\alpha)$ representing) with $\varphi^{G}_{-C_{\psi}(x^{*},\alpha),k}(0)=0$, if $C_{\psi}(x^{*},\alpha)$ is nontrivial and closed, and $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha)$ (respectively, $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$).
  • According to Lemma 2 ($2^{\circ}$, $4^{\circ}$), the Bishop-Phelps type (seminorm-linear) function $\varphi_{x^{*},\alpha}$ introduced in (6) is sublinear and $-C_{\psi}(x^{*},\alpha)$ representing with $\varphi_{x^{*},\alpha}(0)=0$, and if $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, then $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing.
4.1 Basic scalarization results
Theorem 12.
For any $a\in Y$, the following assertions hold:
  • $1^{\circ}$
    If $(x^{*},\alpha)\in K^{a\circ}$, and $\varphi$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing (e.g., if $\varphi$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing), then ${\rm argmin}_{x\in\Omega}\;\varphi(f(x)-a)\subseteq{\rm WEff}(\Omega\mid f,K)$.
  • $2^{\circ}$
    If $(x^{*},\alpha)\in K^{a\#}$ and $\varphi$ is strongly $(C_{\psi}^{>}(x^{*},\alpha)\cup\{0\})$-increasing, then ${\rm argmin}_{x\in\Omega}\;\varphi(f(x)-a)\subseteq{\rm Eff}(\Omega\mid f,K)$.
  • $3^{\circ}$
    Assume that $\overline{x}\in\Omega$. If $(x^{*},\alpha)\in K^{a+}$, $\varphi$ is $C_{\psi}(x^{*},\alpha)$-increasing, and ${\rm argmin}_{x\in\Omega}\;\varphi(f(x)-a)=\{\overline{x}\}$, then $\overline{x}\in{\rm Eff}(\Omega\mid f,K)$.
  • $4^{\circ}$
    If $(x^{*},\alpha)\in K^{a\#}$, $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, and $\varphi$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing or strongly $C_{\psi}(x^{*},\alpha)$-increasing, then ${\rm argmin}_{x\in\Omega}\;\varphi(f(x)-a)\subseteq{\rm PEff}_{He}(\Omega\mid f,K)$.
Proof.
$1^{\circ}$ For $\operatorname{int}\,K=\emptyset$ the result is clear. Assume that $\operatorname{int}\,K\neq\emptyset$. Consider $(x^{*},\alpha)\in K^{a\circ}$. Obviously, if $\varphi$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing, then $\varphi(\cdot-a)$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing as well. Because $(x^{*},\alpha)\in K^{a\circ}$ we have $\emptyset\neq\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$. Finally, Lemma 4 ($2^{\circ}$) and Lemma 5 ($1^{\circ}$) yield the conclusion.
$2^{\circ}$ and $3^{\circ}$ The proof of the assertions is similar to the proof of assertion $1^{\circ}$ by using Lemma 4 ($3^{\circ}$) and Lemma 5 ($2^{\circ},3^{\circ}$).
$4^{\circ}$ Take some $(x^{*},\alpha)\in K^{a\#}$. Since $\varphi$ is strictly (respectively, strongly) $C_{\psi}(x^{*},\alpha)$-increasing we have that $\varphi(\cdot-a)$ is strictly (respectively, strongly) $C_{\psi}(x^{*},\alpha)$-increasing as well. Because $(x^{*},\alpha)\in K^{a\#}$ we get $K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$. Taking into account that $C_{\psi}(x^{*},\alpha)$ is a nontrivial convex cone, we conclude that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$. Finally, with the aid of Lemma 5 ($4^{\circ}$) we get the desired inclusion. ∎
In particular, we like to point out monotonicity properties of the seminorm-linear function $\varphi_{x^{*},\alpha}$ with respect to a Bishop-Phelps type cone $C_{\psi}(x^{*},\alpha)$.
Lemma 13.
Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. Then, the following assertions hold:
  • $1^{\circ}$
    $\varphi_{x^{*},\alpha}$ is $C_{\psi}(x^{*},\alpha)$-increasing.
  • $2^{\circ}$
    If $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, then $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing.
  • $3^{\circ}$
  • $4^{\circ}$
  • $5^{\circ}$
    $\varphi_{x^{*},\alpha}$ is strongly $(C_{\psi}^{>}(x^{*},\alpha)\cup\{0\})$-increasing.
Proof.
$1^{\circ}$ Noting that $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing, the assertion follows from Lemma 4 ($1^{\circ}$).
$2^{\circ}$ If $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$, there is nothing to prove. If $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, using the assumption $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ we have that $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing, and the result follows from Lemma 4 ($1^{\circ}$).
$5^{\circ}$ It follows from the definitions. ∎
The next lemma, which studies the monotonicity properties of $\varphi_{x^{*},\alpha}$ with respect to the original cone $K$, extends a corresponding result derived by Kasimbeyli [8, Th. 3.5] in a normed setting.
Lemma 14.
  • $1^{\circ}$
    $(x^{*},\alpha)\in K^{a+}$ $\iff$ $\varphi_{x^{*},\alpha}$ is $K$-increasing.
  • $2^{\circ}$
  • $3^{\circ}$
Proof.
$1^{\circ}$ By Lemma 13 ($1^{\circ}$), the function $\varphi_{x^{*},\alpha}$ is $C_{\psi}(x^{*},\alpha)$-increasing. If $(x^{*},\alpha)\in K^{a+}$, then $K\subseteq C_{\psi}(x^{*},\alpha)$, hence $\varphi_{x^{*},\alpha}$ is $K$-increasing by Lemma 4 ($3^{\circ}$). Conversely, assuming $\varphi_{x^{*},\alpha}$ is $K$-increasing, for $k\in K$ we have $-k\leq_{K}0$, hence $\varphi_{x^{*},\alpha}(-k)\leq\varphi_{x^{*},\alpha}(0)=0$, which means that $x^{*}(k)-\alpha\psi(k)\geq 0$. Noting that $x^{*}\in K^{+}$, we conclude $(x^{*},\alpha)\in K^{a+}$.
$2^{\circ}$ If $\operatorname{int}\,K=\emptyset$, we have $K^{a\circ}=Y^{*}\times\mathbb{R}_{+}$ and $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing, hence the equivalence is valid. Assume that $\operatorname{int}\,K\neq\emptyset$. If $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, and so $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing by Lemmata 13 ($2^{\circ}$) and 4 ($2^{\circ}$). The proof of the reverse implication is similar to $1^{\circ}$.
$3^{\circ}$ If $(x^{*},\alpha)\in K^{a\#}$, then $K\subseteq C_{\psi}^{>}(x^{*},\alpha)\cup\{0\}$, so $\varphi_{x^{*},\alpha}$ is strongly $K$-increasing by Lemmata 13 ($5^{\circ}$) and 4 ($3^{\circ}$). Also here the proof of the reverse implication is similar to $1^{\circ}$. ∎
Lemma 15.
Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ and assume that $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing. Then, the following assertions are valid:
  • $1^{\circ}$
    If $\bar{x}\in{\rm Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$, then $(\bar{x},0)$ solves the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$ and any $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$. More precisely, $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of ($P_{\varphi}^{a,k}$).
  • $2^{\circ}$
    If $\bar{x}\in{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$, then $(\bar{x},0)$ solves the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$ and any $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
  • $3^{\circ}$
    Consider $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$. If $\{(x,\bar{\lambda})\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem ($P_{\varphi}^{a,k}$) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\rm Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
  • $4^{\circ}$
    If $(\bar{x},\bar{\lambda})$ solves the problem ($P_{\varphi}^{a,k}$) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
Proof.
Follows immediately from Lemma 8 taking into account that $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing. ∎
Theorem 16.
Consider $\bar{x}\in\Omega$. Take some $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, $a\in Y,k\in Y\setminus\{0\}$ and assume that $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing and that $(\bar{x},\bar{\lambda})$ is a solution of the optimization problem ($P_{\varphi}^{a,k}$). Then:
  • $1^{\circ}$
    If $(x^{*},\alpha)\in K^{a\#}$ and $C_{\psi}(x^{*},\alpha)$ is solid, then $\bar{x}\in{\rm PEff}_{He}(\Omega\mid f,K)$.
  • $2^{\circ}$
    If $(x^{*},\alpha)\in K^{a\circ}$, then $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$.
Proof.
$1^{\circ}$ Remark that $\bar{x}\in{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ in view of Lemma 15 ($4^{\circ}$). The latter set is contained in ${\rm PEff}_{He}(\Omega\mid f,K)$ (since $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$).
$2^{\circ}$ For the case $\operatorname{int}\,K=\emptyset$ the statement is clear. Assume that $\operatorname{int}\,K\neq\emptyset$. Again we have $\bar{x}\in{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Here the latter set is contained in ${\rm WEff}(\Omega\mid f,K)$ (since $\emptyset\neq\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$). ∎
Proposition 17.
Let $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ such that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ and $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$
    $\bar{x}\in{\rm Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
  • $2^{\circ}$
  • $3^{\circ}$
    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\rm argmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x}))$.
If there are sets $\mathcal{P}\subseteq Y$, $\mathcal{T}\subseteq C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\rm Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$, then $1^{\circ}-3^{\circ}$ are equivalent to:
  • $5^{\circ}$
    There are $a\in\mathcal{P}$, $k\in\mathcal{T}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))=s$ and $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\rm argmin}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
Proof.
Note that $C_{\psi}(x^{*},\alpha)$ is a convex cone, $\varphi$ is $-C_{\psi}(x^{*},\alpha)$-representing with $\varphi(0)=0$ and $C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))\neq\emptyset$.
$1^{\circ}\iff 2^{\circ}$ The equivalence follows immediately from Lemma 15 ($1^{\circ}$, $3^{\circ}$).
$1^{\circ}\iff 3^{\circ}$ The result follows directly from Lemma 6 ($1^{\circ}$).
$1^{\circ}\iff 4^{\circ}\iff 5^{\circ}$ The result follows from Proposition 9 taking into account that ($P_{\varphi}^{a,k}$) is the same problem as (7) with $C_{\psi}(x^{*},\alpha)$ in the role of $K$. ∎
Proposition 18.
Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ such that $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$
    $\bar{x}\in{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
  • $2^{\circ}$
If $\varphi$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing (e.g., if $\varphi:=\varphi_{x^{*},\alpha}$ and $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$), then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
  • $3^{\circ}$
    $\bar{x}\in{\rm argmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x}))$.
If there are sets $\mathcal{P}\subseteq Y$, $\mathcal{T}\subseteq\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
  • $5^{\circ}$
    There are $a\in\mathcal{P}$, $k\in\mathcal{T}$ and $s\in\mathcal{R}$ so that $s=\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))={\rm min}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
Proof.
First, note that $C_{\psi}(x^{*},\alpha)$ is a solid, convex cone and $\varphi$ is $-C_{\psi}(x^{*},\alpha)$-representing with $\varphi(0)=0$.
$1^{\circ}\iff 2^{\circ}$ The equivalence follows immediately from Lemma 15 ($2^{\circ}$, $4^{\circ}$).
$1^{\circ}\iff 3^{\circ}$ The result follows from Lemma 6 ($2^{\circ}$). Note, since $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ and $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, the function $\varphi$ is strictly $-C_{\psi}(x^{*},\alpha)$-representing by Lemma 2 ($4^{\circ}$) with $\varphi(0)=0$.
$1^{\circ}\iff 4^{\circ}\iff 5^{\circ}$ The result follows directly from the Proposition 10. ∎
Remark 11.
The results derived in Theorem 12 (combined with Lemma 13) and Propositions 17 and 18 for the case $\varphi=\varphi_{x^{*},\alpha}$ extend some known results in the literature, which are given in a normed space $(Y,\psi)$, $\psi(\cdot):=||\cdot||$, preordered by a convex cone $K\subseteq Y$:
  • Gasimov [6, Th. 1] proved a corresponding statement as in Theorem 12 ($4^{\circ}$) for Benson proper efficiency (instead of Henig proper efficiency).
  • Kasimbeyli [9, Th. 5.4] stated similar assertions as in Theorem 12 for a closed, pointed, convex cone $K$ and Benson proper efficiency (instead of Henig proper efficiency).
  • Gasimov [6, Th. 2] (see also Guo and Zhang [44, Th. 4.5]) stated that $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,C_{||\cdot||}(x^{*},\alpha))$ (with $\mathbb{A}$ considered in Benson proper efficiency) implies $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\rm argmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
Remark 12.
Assume that $K$ is a nontrivial cone and $(x^{*},\alpha)\in K^{a\#}$. Then, $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)\subseteq C_{\psi}(x^{*},\alpha)\setminus\{0\}$, hence $\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\rm Eff}(\Omega\mid f,K)$. If further $K$ has the normlike-base $B_{K}(\psi)$, and $C_{\psi}(x^{*},\alpha)$ is solid (e.g. if $K$ is solid or $\psi$ is continuous), then $C_{\psi}(x^{*},\alpha)$ is a nontrivial, convex cone with $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, which means that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$, and so $\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\rm PEff}_{He}(\Omega\mid f,K)$.
If $K$ is a nontrivial, solid cone and $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, hence $\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\rm WEff}(\Omega\mid f,K)$.
4.2 Nonlinear Cone Separation Conditions